# Some lemmas on spectral radius of graphs: including an application

### v1

@@ Abstract
For a graph $G$, the spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. In this paper, we give three lammas on $\rho(G)$ when $G$ contains a spanning complete bipartite graph. Moreover, an application was also included at the end.
Keywords: spectral radius; walk; cycle; planar graph.
2020 Mathematics Subject Classification: 05C50.
@@ 1 Introduction
Let $G$ be a graph with vertices $v_{1},v_{2},...,v_{n}$. The adjacency matrix of $G$ is an $n\times n$ matrix $(a_{ij})$, where $a_{ij}=1$ if $v_{i}\sim v_{j}$, and $a_{ij}=0$ otherwise. The spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. By the well-known Perron–Frobenius theorem, $\rho(G)$ has a non-negative eigenvector. A non-negative eigenvector corresponding to $\rho(G)$ is called a Perron vector of $G$. If $G$ is connected, any Perron vector of $G$ has positive entries. The spectral radius of a graph is closely related to its convergence property. In this paper, we give the following three lammas on spectral radius of graphs, which will be used.
@@
For integers $n_{1}\geq n_{2}+2\geq 1$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|\geq 0$. Then $\rho(G_{1})>\rho(G_{2})$.
@@
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq 1$ and $n_{1}\geq n_{2}+2$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 4}n_{s}\geq 100$. Then $\rho(G_{1})<\rho(G_{2})$.
@@
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq n_{5}\geq 1$ and $n_{1}\geq n_{2}+3$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 5}n_{s}\geq 220$. Then $\rho(G_{1})<\rho(G_{2})$.
The rest of the paper is organized as follows. In Section 2, we include several lemmas, which will be used later. The proofs of Lemmas 1.1, 1.2 and 1.3 are given in Section 3. In Section, we include an application on spectral radius of $C_{\ell}$-free planar graphs of order $n\geq 1.8\times 10^{17}$, where $5\leq\ell\leq n$.
@@ 2 Preliminaries
Let $G$ be a graph. For a vertex $u$ and a walk $Q$ of $G$, we say that $Q$ crosses $u$, if $u$ is one of the vertices in the sequence of $Q$. For any integer $\ell\geq 1$ and $u,v\in V(G)$, let $W^{\ell}_{u,v}(G)$ be the number of walks $Q$ of length $\ell$ in $G$, such that $Q$ crosses $u$ and $v$. Let $P_{0}$ denote the empty graph.
@@
Let $G$ be a path of order $n\geq 3$. Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$. Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
@@
(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $a\geq 0$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”). Thus,
$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=2\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
@@
$\leq\frac{8}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
$=\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
$\leq\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
$\leq\frac{8}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
$=\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
@@
$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
@@
$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}<e^{0.52}<1.7$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{4\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$
@@
Now consider $g(x)=5.8+\frac{6.8}{x-1}+\frac{16}{x-4}-x$ for $x\geq 10$. Clearly, $g(x)$ is decreasing for $x\geq 10$. By a calculation, $g(10)=-\frac{7}{9}<0$. Thus $g(x)<0$ for any $x\geq 10$. Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{220},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}<e^{0.405}<1.5$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$

### v2

@@ Abstract
For a graph $G$, the spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. In this paper, we give three lammas on $\rho(G)$ when $G$ contains a spanning complete bipartite graph. Using these lemmas and typical spectral method, we characterized the unique extremal graph with the maximum spectral radius among all planar graphs of large order $n$ without a cycle of length $\ell$, where $5\leq\ell\leq n$.
Keywords: spectral radius; planar graph; cycle; walk.
2020 Mathematics Subject Classification: 05C50; 05C35.
@@ 1 Introduction
Let $G$ be a graph with vertices $v_{1},v_{2},...,v_{n}$. The adjacency matrix $A(G)$ of $G$ is an $n\times n$ matrix $(a_{ij})$, where $a_{ij}=1$ if $v_{i}\sim v_{j}$, and $a_{ij}=0$ otherwise. The spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. By the well-known Perron–Frobenius theorem, $\rho(G)$ has a non-negative eigenvector. A non-negative eigenvector corresponding to $\rho(G)$ is called a Perron vector of $G$. If $G$ is connected, any Perron vector of $G$ has positive entries. The spectral radius of a graph is closely related to its convergence property. In this paper, we first give the following three lammas on spectral radius of graphs. (Let $P_{0}$ denote the empty graph.)
@@
For integers $n_{1}\geq n_{2}+2\geq 3$, let
$G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup T)$
and
$G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup T),$
where $H,T$ are two graphs with $|H|\geq 1$ and $|T|\geq 0$. Then $\rho(G_{1})>\rho(G_{2})$.
@@
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq 1$ and $n_{1}\geq n_{2}+2$, let
$G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T)$
and
$G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T),$
where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 4}n_{s}\geq 130$. Then $\rho(G_{1})<\rho(G_{2})$.
@@
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq n_{5}\geq 1$ and $n_{1}\geq n_{2}+3$, let
$G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T)$
and
$G_{2}=H\vee(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T),$
where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 5}n_{s}\geq 310$. Then $\rho(G_{1})<\rho(G_{2})$.
A graph is called planar, if it can be drawn in the plane without crossed edges. Tait and Tobin [6] proved that $K_{2}\vee P_{n-2}$ is the only extremal graph with the maximum spectral radius among all planar graphs of large order $n$. This (asymptotically) solves a conjecture proposed by Boots and Royle [1], and independently by Cao and Vince [3]. Recently, the researchers are also interested in determining the extremal graphs with the maximum spectral radius among all $C_{\ell}$-free planar graphs of large order $n$, where $3\leq\ell\leq n$. The case $\ell=4$ can be settled from the results of Nikiforov [5], and Zhai and Wang [8]. Very recently, Fang, Lin and Shi [4] solved the case when $\ell=3$ and $5\leq\ell\ll n$. Xu, Lin and Fang [7] solved the case when $\frac{\ell}{n}$ is close to 1. They [7] proposed the general case as a problem. Armed with Lemma 1.1 and Lemma 1.2, using typical spectral method we can solve this problem. That is, we have the following Theorem 1.4.
For a finite family of graphs $\mathcal{G}$, let ${\rm SPEX}(\mathcal{G})$ be the set of all extremal graphs with the maximum spectral radius in $\mathcal{G}$.
Theorem 1.4
The rest of the paper is organized as follows. In Section 2, we include several lemmas, which will be used later. The proofs of Lemmas 1.1, 1.2 and 1.3 are given in Section 3. In Section 4, we give the proof of Theorem 1.4.
@@ 2 Preliminaries
Let $G$ be a graph. For a vertex $u$ and a walk $Q$ of $G$, we say that $Q$ crosses $u$, if $u$ is one of the vertices in the sequence of $Q$. For any integer $\ell\geq 1$ and $u,v\in V(G)$, let $W^{\ell}_{u,v}(G)$ be the number of walks $Q$ of length $\ell$ in $G$, such that $Q$ crosses $u$ and $v$. Recall that $P_{0}$ denotes the empty graph.
@@
The following fact can be proved by induction on $\ell$.
Fact 1. For integers $n\geq 4\ell\geq 0$, we have
$\binom{n}{0}+\binom{n}{1}+\cdots+\binom{n}{\ell}\leq\frac{5}{4}\frac{n^{\ell}}{\ell!}$
@@
Let $G$ be a path of order $n\geq 3$. Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$. Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
@@
(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $0\leq a\leq\ell$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”). Thus, using Fact 1 as $\ell+a\geq 4\lfloor\frac{a-i}{2}\rfloor$,
$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{0}+\binom{\ell+a}{1}+\cdots+\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{5}{4}\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{5}{4}\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times\frac{5}{4}\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=\frac{5}{2}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
@@
$\leq\frac{10}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
$=\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
$\leq\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
$\leq\frac{10}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
$=\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
@@
$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
@@
$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{4}+1)+1}{\rho_{1}^{2}}}\leq e^{0.51}<1.67$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{5\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$
@@
Now consider $g(x)=8.5+\frac{7.5}{x-1}+\frac{16}{x-4}-x$ for $x\geq\sqrt{130}$. Clearly, $g(x)$ is decreasing for $x\geq\sqrt{130}$. By a calculation, $g(\sqrt{130})=-0.01\cdots<0$. Thus $g(x)<0$ for any $x\geq\sqrt{130}$. Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
@@
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
@@
$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{310},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{5}+\frac{6}{5})+\frac{4}{5}}{\rho_{1}^{2}}}<e^{0.403}<1.5$. Thus,
@@
$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$